Lily is 13 years older than Richie. Next year, she will be twice as old as he is. How old are both of them of them now?
This is I have...
L+13=R



Answer by stanbon(75887)

Lily is 13 years older than Richie. Next year, she will be twice as old as he is. How old are both of them of them now?
------------
Equations:
NoW DATA:
Let Lily be "L" years old
Let Rickie be "R" years old
---
Next Year DATA"
Lily will be L+1
Rickie will be R + 1
-----------------------
Equations:
L = R + 13
L+1 = 2(R + 1)
---
Substitue for "L" and solve for R:
R+13+1 = 2R + 2
R + 14 = 2R + 2
R = 12 (Rickie's age now)
---
Since L = R + 13, L = 12+13 = 25 (Lily's age now)
=================================================
Cheers,
Stan H.